Adam is willing to pay $100 for his 1st pack of toilet paper, $25 for his 2nd, and $10 for his 3rd. Bella is willing to pay $80 for her 1st pack of toilet paper, $70 for her 2nd, and $15 for her 3rd. Chas is willing to pay $30 each for as many packs of toilet paper as he can get. The store currently only has 3 packs left, and only these 3 customers will be at the store to buy toilet paper. Calculate overall consumer surplus for each of the following scenarios: a. The store sells the toilet paper for $25/pack, but Chas gets there first and buys all 3 packs. b. The store keeps the price at $25/pack, but limits purchases to 1 pack per customer. c. The store raises the price to $35/pack.
Answer #1

10 (a) Consumer Surplus = Maximum willingnes to pay - Actual price Y the store sells the toilet paper for $25 / park & Chas5 (100-25) + (80-25) + (30-25) = 75 +55+5 = $135 i( 4 the store raises the brice to $35 perck then chas will not burchase any toilet paper as his willingness to pay is less th